# Digging into the ultraviolet luminosity functions of galaxies at high redshifts: galaxies evolution, reionization, and cosmological parameters

### v1

@@ 3.2 Properties of the Reionization
$z_{0}$ Uniform(0,10) $5.38^{+0.65}_{-0.70}\,(5.52^{+0.60}_{-0.79})$
$z_{\rm max}$ Uniform(10,30) $20.58^{+6.25}_{-6.75}\,(20.33^{+6.20}_{-6.69})$
$f_{\rm esc}$b Uniform(0,1) $0.43^{+0.02}_{-0.02}\,(0.28^{+0.01}_{-0.02})$
@@
    a The values in these rows represent $\rho_{\rm UV}$ follows $M_{\rm trunc}=-15$. The uncertainties correspond to the $68\%$ confidence level.
@@
    b During fitting process, $\xi_{\rm ion}$ is assumed to be $10^{24.5}$ for convenient analysis, but $\langle f_{\rm esc}\xi_{\rm ion}\rangle$ is still the practical fitting parameters.
Whereafter, the evolution of $\mathcal{Q}_{\rm H_{II}}$ can be constrained by the observations of the neutral hydrogen fraction $x_{\rm H_{I}}$ ($x_{\rm H_{I}}=1-\mathcal{Q}_{\rm H_{II}}$). These additional data were obtained through various methods, including analysis of the ${\rm Ly}\alpha$ and ${\rm Ly}\beta$ forests of quasars, distributions of ${\rm Ly}\alpha$ equivalent width, adsorptions in ${\rm Ly}\alpha$ damping wings, fractions of ${\rm Ly}\alpha$ emitters, the afterglow spectrum of the gamma-ray burst, and the Gunn Peterson troughs. All of the observation results are summarized in Table 2.
@@
where $\sigma_{\rm T}$ is the Thomson cross-section ($\sigma_{\rm T}=6.65\times 10^{-29}\,{\rm m^{2}}$), $\eta=1$ at $z>4$ and $\eta=2$ at $z\leq 4$. Thus, the observation of $\tau_{\rm e}$ ($\tau_{\rm e}=0.054\pm 0.007$) can be used to constrain $\mathcal{Q}_{\rm H_{II}}$. The Bayesian analysis follows the same process described in subsection 2.4. For these upper and lower limits, the likelihood functions are assumed to be uniform within the limited range and to be a half-Gaussian beyond the boundary (Greig & Mesinger, 2017). The prior distributions and the posterior results of the free parameters are listed in Table 3. Using the $\rho_{\rm UV}$ derived from UV LFs (at $z>10$, the $\rho_{\rm UV}$ is extrapolated), the estimated evolution of $x_{\rm H_{I}}$ is shown in Figure 4. Furthermore, we find out that the reionization started at $20.58^{+6.25}_{-6.75}\,(20.33^{+6.20}_{-6.69})$ and ended at $5.38^{+0.65}_{-0.70}\,(5.52^{+0.60}_{-0.79})$ when $M_{\rm trunc}$ is fixed to $-17\,(-15)$. Interestingly, such results are insensitive on the choice of $M_{\rm trunc}$.
Since Equation 17 describes the relation between $\mathcal{Q}_{\rm H_{II}}$ and $\tau_{\rm e}$, we evaluate $\tau$ by fitting observations of $x_{\rm H_{I}}$ listed in Table 2. Using $\rho_{\rm UV}$ with $M_{\rm trunc}=-17$, we present the extrapolated projections of $\tau_{\rm e}$ in Figure 5. Notably, our estimation is consistent with the result of Planck Collaboration et al. (2020), demonstrating the validity and reliability of our model.
@@ 4 Conclusions and discussions
Under the framework of $\Lambda$CDM, the UV LFs within the redshift range of $z=4-10$ can constrain the evolution of SFE stringently. In Figure 2, the profile of SFE shows a clear tendency of evolution with redshift, particularly in the low mass range. The corresponding mass of the DM halos ($\sim 10^{12}\,\rm M_{\odot}$) at maximum SFE ($\sim 20\%$) presents a shift towards higher values, although further investigation is needed/warranted. Since the variable feedback mechanism and the environmental factors influence the relative strength of SFE, disentangling the contributions of each component solely from UV LFs poses challenges. Fortunately, dynamical simulations provide comparisons for these elements(e.g., Guszejnov et al., 2022; Scharré et al., 2024), including the IMF, stellar radiation, stellar winds, supernovae, AGN and others. Therefore, once the profile of SFE can be constrained by other methods beyond UV LFs, the cosmological researches relying on UV LFs can yield more precise results, as the intrinsic degeneracy between SFE and cosmological parameters is alleviated. Furthermore, based on the derived UV luminosity density and abundant observations of the IGM neutral fraction, the beginning and ending redshifts of the reionization epoch are constrained to $20.58^{+6.25}_{-6.75}$ ($20.33^{+6.20}_{-6.69}$) and $5.38^{+0.65}_{-0.70}$ ($5.52^{+0.60}_{-0.79}$) with $M_{\rm trunc}=-17$ and $-15$, respectively. If considering the observations of UV LFs at $z>10$, $\rho_{\rm UV}$ will be much higher than the extrapolated ones (Donnan et al., 2023; Harikane et al., 2023), but does not impact Figure 4 notably. We validate the reliability of our model by comparing the inferred Thomson scattering optical depth $\tau_{\rm e}$ with the result of Planck Collaboration et al. (2020). Nonetheless, our model is applicable only to the nonlinear regime and not to regimes with $z>10$, as uncertainties in stellar populations and dust attenuation lead to large uncertainties of SFE (Inayoshi et al., 2022; Wang et al., 2023) and other parameters.
@@
On the other hand, we attempt to analyze cosmology by UV LF observations. Similar to previous works, we find that only introducing some reasonable and additional information, the cosmological parameters can be constrained within physically meaningful scopes. Within a specific parameter space, we obtain $\sigma_{8}=0.815^{+0.026}_{-0.030}$, which is consistent with the $\Lambda$CDM framework. Moreover, following Sabti et al. (2022b) and employing the same dataset and parameter spaces, the inferred $\sigma_{8}=0.784^{+0.051}_{-0.048}$ has a better precision, improved by $\sim 2-3$ times compared to the results reported by Sabti et al. (2022b). It is in line with expectations since $\sigma_{8}$ is constrained by UV LFs mainly and the number of UV LF observations we used is over four times greater than they used, thereby reducing the error to $\sim 0.12/\sqrt{4}$. Besides the increase of data points, the expansion of galaxy survey areas decrease the uncertainties of cosmic variance and then improve the accuracy of cosmology analysis.
@@ 5 Acknowledgements
This work is supported in part by NSFC under grants of No. 11921003, No. 12233011 and 12303056; S.-P. T. acknowledges support from the General Fund (No. 2023M733736) of the China Postdoctoral Science Foundation; G.-W. Y. acknowledges support from the University of Trento and the Provincia Autonoma di Trento (PAT, Autonomous Province of Trento) through the UniTrento Internal Call for Research 2023 grant “Searching for Dark Energy off the beaten track” (DARKTRACK, grant agreement no.E63C22000500003, PI: Sunny Vagnozzi).
Software: Pymultinest (Buchner (2016), version 2.11, https://pypi.org/project/pymultinest/), HMFcalc (Murray (2014)), https://github.com/halomod/hmf/
Appendix A The posterior distributions of the parameters of SFE model

### v2

@@ 3.2 Properties of the Reionization
$z_{0}$ Uniform(0,10) $5.3^{+0.8}_{-1.0}\,(5.3^{+0.8}_{-1.0})$
$z_{\rm max}$ Uniform(10,30) $18.8^{+7.2}_{-6.0}\,(19.0^{+6.7}_{-5.8})$
$\langle f_{\rm esc}\xi_{\rm ion}\rangle$b Uniform(0,1) ×$\times 10^{25.5}$ $0.3^{+0.1}_{-0.1}\,(0.4^{+0.1}_{-0.1})\times 10^{25.5}$
@@
    a The values in these rows represent $\rho_{\rm UV}$ follows $M_{\rm trunc}=-17$. The uncertainties correspond to the $68\%$ confidence level.
@@
    b Assuming $\xi_{\rm ion}=10^{25.6}\,(10^{25.8})$ (Tang et al., 2023; Atek et al., 2024), the values of $f_{\rm esc}$ is $\sim 20\%$.
Whereafter, the evolution of $\mathcal{Q}_{\rm H_{II}}$ can be constrained by the observations of the neutral hydrogen fraction $x_{\rm H_{I}}$ ($x_{\rm H_{I}}=1-\mathcal{Q}_{\rm H_{II}}$). These additional data were obtained through various methods, including analysis of the ${\rm Ly}\alpha$ and ${\rm Ly}\beta$ forests of quasars, distributions of ${\rm Ly}\alpha$ equivalent width, adsorptions in ${\rm Ly}\alpha$ damping wings, fractions of ${\rm Ly}\alpha$ emitters, the afterglow spectrum of the gamma-ray burst, and the Gunn Peterson troughs. Using the similar screening methodology of UV LF observations, each of the data points of $x_{\rm H_{I}}$ at the same redshift has been confirmed without any redundancy. All of the observation results are summarized in Table 2.
@@
where $\sigma_{\rm T}$ is the Thomson cross-section ($\sigma_{\rm T}=6.65\times 10^{-29}\,{\rm m^{2}}$), $\eta=1$ at $z>4$ and $\eta=2$ at $z\leq 4$. Thus, the observation of $\tau_{\rm e}$ ($\tau_{\rm e}=0.054\pm 0.007$ from Planck Collaboration et al. (2020)) can be used to constrain $\mathcal{Q}_{\rm H_{II}}$. The Bayesian analysis follows the same process described in subsection 2.5. For these upper and lower limits, the likelihood functions are assumed to be uniform within the limited range and to be a half-Gaussian beyond the boundary (Greig & Mesinger, 2017). The prior distributions and the posterior results of the free parameters are listed in Table 3. Using the $\rho_{\rm UV}$ derived from UV LFs (at $z>10$, the $\rho_{\rm UV}$ is extrapolated), the estimated evolution of $x_{\rm H_{I}}$ is shown in Figure 5. Furthermore, we find out that the reionization started at $18.8^{+7.2}_{-6.0}\,(19.0^{+6.7}_{-5.8})$ and ended at $5.3^{+0.8}_{-1.0}\,(5.3^{+0.8}_{-1.0})$ when $M_{\rm trunc}$ is fixed to $-15\,(-17)$. Interestingly, such results are insensitive on the choice of $M_{\rm trunc}$.
Since Equation 17 describes the relation between $\mathcal{Q}_{\rm H_{II}}$ and $\tau_{\rm e}$, we evaluate $\tau$ by fitting observations of $x_{\rm H_{I}}$ listed in Table 2. Using $\rho_{\rm UV}$ with $M_{\rm trunc}=-15$, we present the extrapolated projections of $\tau_{\rm e}$ in Figure 6. Notably, our estimation is consistent with the result of Planck Collaboration et al. (2020), demonstrating the validity and reliability of our model.
@@ 4 Conclusions and discussions
Under the framework of $\Lambda$CDM, the UV LFs within the redshift range of $z=4-10$ can constrain the evolution of SFE stringently. In Figure 2, the profile of SFE shows a clear tendency of evolution with redshift, particularly in the low mass range. The corresponding mass of the DM halos ($\sim 10^{12}\,\rm M_{\odot}$) at maximum SFE ($\sim 20\%$) presents a shift towards higher values, although further investigation is needed/warranted. Since the variable feedback mechanism and the environmental factors influence the relative strength of SFE, disentangling the contributions of each component solely from UV LFs poses challenges. Fortunately, dynamical simulations provide comparisons for these elements(e.g., Guszejnov et al., 2022; Scharré et al., 2024), including the IMF, stellar radiation, stellar winds, supernovae, AGN and others. Therefore, once the profile of SFE can be constrained by other methods beyond UV LFs, the cosmological researches relying on UV LFs can yield more precise results, as the intrinsic degeneracy between SFE and cosmological parameters is alleviated. Furthermore, based on the derived UV luminosity density and abundant observations of the IGM neutral fraction, the beginning and ending redshifts of the reionization epoch are constrained to $18.8^{+7.2}_{-6.0}\,(19.0^{+6.7}_{-5.8})$ and $5.3^{+0.8}_{-1.0}\,(5.3^{+0.8}_{-0.9})$ with $M_{\rm trunc}=-17$ and $-15$, respectively. If considering the observations of UV LFs at $z>10$, $\rho_{\rm UV}$ will be much higher than the extrapolated ones (Donnan et al., 2023; Harikane et al., 2023), but does not impact Figure 5 notably. We validate the reliability of our model by comparing the inferred Thomson scattering optical depth $\tau_{\rm e}$ with the result of Planck Collaboration et al. (2020). Nonetheless, our model is applicable only to the nonlinear regime and not to regimes with $z>10$, as uncertainties in stellar populations and dust attenuation lead to large uncertainties of SFE (Inayoshi et al., 2022; Wang et al., 2023) and other parameters.
@@
On the other hand, we attempt to analyze cosmology by UV LF observations. Similar to previous works, we find that only introducing some reasonable and additional information, the cosmological parameters can be constrained within physically meaningful scopes. Within a specific parameter space, we obtain $\sigma_{8}=0.816^{+0.029}_{-0.033}$, which is consistent with the $\Lambda$CDM framework. Moreover, following Sabti et al. (2022b) and employing the same dataset and parameter spaces, the inferred $\sigma_{8}=0.796^{+0.056}_{-0.053}$ has a better precision, improved by $\sim 2-3$ times compared to the results reported by Sabti et al. (2022b). It is in line with expectations since $\sigma_{8}$ is constrained by UV LFs mainly and the number of UV LF observations we used is over four times greater in number than they used, thereby reducing the error to $\sim 0.12/\sqrt{4}$. In addition to increasing the number of data points, expanding the survey area reduces the uncertainties associated with cosmic variance, thereby enhancing the accuracy of parameter estimation.
@@ 5 Acknowledgements
We thank the anonymous referee for helpful comments and suggestions. This work is supported in part by NSFC under grants of No. 11921003, No. 12233011 and 12303056; S.-P. T. acknowledges support from the General Fund (No. 2023M733736) of the China Postdoctoral Science Foundation and the Postdoctoral Fellowship Program of CPSF (GZB20230839); G.-W. Y. acknowledges support from the University of Trento and the Provincia Autonoma di Trento (PAT, Autonomous Province of Trento) through the UniTrento Internal Call for Research 2023 grant “Searching for Dark Energy off the beaten track” (DARKTRACK, grant agreement no.E63C22000500003, PI: Sunny Vagnozzi).
Software: Nessai (Williams et al. (2021)), version 0.12.0, https://nessai.readthedocs.io/en/latest/, HMFcalc (Murray (2014)), https://github.com/halomod/hmf/
Appendix A The contributions of cosmic variance
In this section, we present the impact of cosmic variance across different magnitudes and redshift. As an example, we list the observations from Bouwens et al. (2021). The total uncertainties used in our analysis include both the observed uncertainties and the contributions from cosmic variance.
{ruledtabular}
Table 4: The rest-frame UV LFs (Bouwens et al., 2021) and the corresponding cosmic variance (CV) at $z\sim 4-9$
Area $M_{\rm UV}$ $\Phi$ CV Area $M_{\rm UV}$ $\Phi$ CV
($\rm{arcmin}^{2}$) ($\rm{mag}^{-1}\rm{Mpc}^{-3}$) ($\rm{arcmin}^{2}$) ($\rm{mag}^{-1}\rm{Mpc}^{-3}$)
-22.69 $0.000005\pm 0.000004$ $0.2871$a
-22.19 $0.000015\pm 0.000009$ $0.2871$22.19 $0.000001\pm 0.000002$ $0.2079$
-21.69 $0.000144\pm 0.000022$ $0.2283$21.69 $0.000041\pm 0.000011$ $0.1657$
-21.19 $0.000344\pm 0.000038$ $0.1847$21.19 $0.000047\pm 0.000015$ $0.1403$
-20.69 $0.000698\pm 0.000068$ $0.1525$20.69 $0.000198\pm 0.000036$ $0.1219$
311.7 -20.19 $0.001624\pm 0.000131$ $0.1292$ 770.2 -20.19 $0.000283\pm 0.000066$ $0.1075$
$z\sim 4$19.69 $0.002276\pm 0.000199$ $0.1109$ $z\sim 7$19.69 $0.000589\pm 0.000126$ $0.0958$
-19.19 $0.003056\pm 0.000388$ $0.0984$19.19 $0.001172\pm 0.000336$ $0.0862$
-18.69 $0.004371\pm 0.000689$ $0.0882$18.69 $0.001433\pm 0.000419$ $0.0785$
-17.94 $0.010160\pm 0.000920$ $0.0760$17.94 $0.005760\pm 0.001440$ $0.0697$
-16.94 $0.027420\pm 0.003440$ $0.0638$16.94 $0.008320\pm 0.002900$ $0.0612$
-15.94 $0.028820\pm 0.008740$ $0.0547$
-23.11 $0.000001\pm 0.000001$ $0.2989$
-22.61 $0.000004\pm 0.000002$ $0.2305$
-22.11 $0.000028\pm 0.000007$ $0.1814$21.85 $0.000003\pm 0.000002$ $0.1603$
-21.61 $0.000092\pm 0.000013$ $0.1456$21.35 $0.000012\pm 0.000004$ $0.1395$
-21.11 $0.000262\pm 0.000024$ $0.1192$20.85 $0.000041\pm 0.000011$ $0.1234$
770.2 -20.61 $0.000584\pm 0.000044$ $0.1006$ 988.5 -20.10 $0.000120\pm 0.000040$ $0.1047$
$z\sim 5$20.11 $0.000879\pm 0.000067$ $0.0865$ $z\sim 8$19.35 $0.000657\pm 0.000233$ $0.0902$
-19.61 $0.001594\pm 0.000156$ $0.0763$18.60 $0.001100\pm 0.000340$ $0.0799$
-19.11 $0.002159\pm 0.000346$ $0.0683$17.60 $0.003020\pm 0.001140$ $0.0696$
-18.36 $0.004620\pm 0.000520$ $0.0587$
-17.36 $0.008780\pm 0.001540$ $0.0496$
-16.36 $0.025120\pm 0.007340$ $0.0425$
-22.52 $0.000002\pm 0.000002$ $0.2423$
-22.02 $0.000014\pm 0.000005$ $0.1892$
-21.52 $0.000051\pm 0.000011$ $0.1511$21.92 $0.000001\pm 0.000001$ $0.1457$
-21.02 $0.000169\pm 0.000024$ $0.1251$21.12 $0.000007\pm 0.000003$ $0.1273$
770.2 -20.52 $0.000317\pm 0.000041$ $0.1063$ 918.2 -20.32 $0.000026\pm 0.000009$ $0.1124$
$z\sim 6$20.02 $0.000724\pm 0.000087$ $0.0929$ $z\sim 9$19.12 $0.000187\pm 0.000150$ $0.0949$
-19.52 $0.001147\pm 0.000157$ $0.0824$17.92 $0.000923\pm 0.000501$ $0.0814$
-18.77 $0.002820\pm 0.000440$ $0.0709$
-17.77 $0.008360\pm 0.001660$ $0.0598$
-16.77 $0.017100\pm 0.005260$ $0.0518$
  • a
    a Because this data point exceeds the limitation of GALCV, the value of CV is assumed to be the latter one.
Appendix B The posterior distributions of the parameters of SFE model
@@
Appendix C The comparison between our results and GALLUMI
In Figure 11, both the results are assumed to be within the $\Lambda CDM$ framework. Our findings provide tighter constraints on $\dot{M}_{*}/\dot{M}_{\rm h}$ due to the larger and more precise datasets used. The minor differences between the two results primarily stem from the variations in the observations. Additionally, a slight discrepancy in the definition of $\dot{M_{\rm h}}$ between the two models may also contribute to this mild discrepancy.
Figure 11: The posterior distribution of $\dot{M}_{*}/\dot{M}_{\rm h}$. The GALLUMI results use the same HST observations as Sabti et al. (2022b). The solid lines indicate the best-fit values, while the colored regions represent the $90\%$ credible intervals, corresponding to the observed magnitude range.